OpenStudy (trojanpoem):

Prove that the following D.E is exact equation, find the solution: dy = (y/x - cosec^2 (y/x)) dx I know it's homogeneous but that's the question.

OpenStudy (faiqraees):

what does he mean by exact equation?

OpenStudy (welshfella):

\[(\delta f / \delta x) * dx + (\delta f / \delta y) * dy = 0\]

OpenStudy (welshfella):

for an exact equation

OpenStudy (welshfella):

an example of an exact equation would be xy' + y = e^x

OpenStudy (welshfella):

this can be rearranged to x dy + (y - e^x) dx = 0 we see that the partial derivatives are both = 1 so the equation is exact.

OpenStudy (phi):

If we let u= y/x and thus u x = y from which we get dy= u dx + x du your equation can be written as \[ u \ dx + x \ du = u - \csc^2 u \ dx \\ x \ du = -\csc^2 u \ dx \] or \[ x\ du = - \frac{1}{\sin^2 u}\ du \\ \frac{1}{x} \ dx +\sin^2 u \ du =0\] the equation is exact if \[ \frac{\partial}{\partial u} \frac{1}{x} =\frac{\partial}{\partial x} \sin^2 u\] which it is, 0=0

OpenStudy (holsteremission):

\[\mathrm dy = \left(\frac{y}{x}-\csc^2\frac{y}{x}\right)\,\mathrm dx\iff \mathrm dy+\left(\csc^2\frac{y}{x}-\frac{y}{x}\right)\,\mathrm dx=0\]The ODE is exact if \[\frac{\partial}{\partial x}1=\frac{\partial}{\partial y}\left[\csc^2\frac{y}{x}-\frac{y}{x}\right]\]The LHS is clearly \(0\), so for the ODE to be exact this must also be the case for the RHS. You have \[\frac{\partial}{\partial y}\left[\csc^2\frac{y}{x}-\frac{y}{x}\right]=-\frac{1}{x}\left(1+2\csc^2\frac{y}{x}\cot\frac{y}{x}\right)\neq0\]so the equation, as is, is *not* exact, but can be made to be with a substitution, as @phi has shown.

OpenStudy (trojanpoem):

@phi, the best. @HolsterEmission, Thanks too bro ^^